OpenStudy (anonymous):

How many solutions does the equation 4p + 7 = 3 + 4 + 4p have?

OpenStudy (anonymous):

help I will medal and fan

OpenStudy (texaschic101):

4p + 7 = 3 + 4 + 4p --- combine like terms 4p + 7 = 7 + 4p if they are the same...then there is infinite solutions

OpenStudy (anonymous):

thank you

OpenStudy (texaschic101):

YW :)

OpenStudy (anonymous):

I need a help on other questions

OpenStudy (texaschic101):

ok....ask away